# Automating eHMI Action Design with LLMs for Automated Vehicle Communication

### v1

@@ Abstract
The absence of explicit communication channels between automated vehicles (AVs) and other road users requires the use of external Human-Machine Interfaces (eHMIs) to convey messages effectively in uncertain scenarios. Currently, most eHMI studies employ predefined text messages and manually designed actions to convey these messages, which limits the real-world deployment of eHMIs, where adaptability in dynamic scenarios is essential. Given the generalizability and versatility of large language models (LLMs), they could potentially serve as automated action designers for the message-action design task. To validate this idea, we make three contributions: (1) We propose a pipeline that integrates LLMs and 3D renderers, using LLMs as action designers to generate executable actions for controlling eHMIs and rendering action clips. (2) We collect a user-rated Action-Design Scoring dataset comprising a total of 320 action sequences for eight intended messages and four representative eHMI modalities. The dataset validates that LLMs can translate intended messages into actions close to a human level, particularly for reasoning-enabled LLMs. (3) We introduce two automated raters, Action Reference Score (ARS) and Vision-Language Models (VLMs), to benchmark 18 LLMs, finding that the VLM aligns with human preferences yet varies across eHMI modalities.  [^1] The source code, prompts, Blender scenarios, and rendered clips are available at https://anonymous.4open.science/r/eHMI_action_design/
@@
Ding Xia1, Xinyue Gui1, Fan Gao1, Dongyuan Li1 Mark Colley1,2, Takeo Igarashi1, 1The University of Tokyo, 2University College London,
@@ 2.1 Rule-Based eHMI Action Planning
Currently, eHMI action planning generally follows a fixed design approach, in which human designers establish behavioral rules based on the specific features of different eHMI modalities. For example, in text- and icon-based eHMIs, designers create content referencing traffic regulation signs or standard messages Eisele and Petzoldt (2022); Eisma et al. (2021). In color- and light-band-based eHMIs, they design the content relying on human intuitive empathy and empirical evaluation with colors and blinking frequencies Bazilinskyy et al. (2019); Dey et al. (2020b). For anthropomorphic eHMIs, such as eyes or arms, designers mimic nonverbal communication cues drawn from common human-human interactions Mahadevan et al. (2018); Ochiai and Toyoshima (2011). Most recently, Colley et al. (2025) proposes using Human-In-The-Loop Multi-Objective Bayesian Optimization to create appropriate eHMIs.
@@ 3.2 Action-Design Scoring Dataset
Robotic eyes are mounted on the front of the autonomous vehicle. The pupil’s position is specified using polar coordinates: the angle spans $[0^{\circ},360^{\circ}]$ (starting from “up” and moving counterclockwise), and the distance spans $[0,1]$, where 00 denotes the center and $1$ is the edge Chang et al. (2022); Gui et al. (2022).

### v2

@@ Abstract
The absence of explicit communication channels between automated vehicles (AVs) and other road users requires the use of external Human-Machine Interfaces (eHMIs) to convey messages effectively in uncertain scenarios. Currently, most eHMI studies employ predefined text messages and manually designed actions to convey these messages, which limits the real-world deployment of eHMIs, where adaptability in dynamic scenarios is essential. Given the generalizability and versatility of large language models (LLMs), they could potentially serve as automated action designers for the message-action design task. To validate this idea, we make three contributions: (1) We propose a pipeline that integrates LLMs and 3D renderers, using LLMs as action designers to generate executable actions for controlling eHMIs and rendering action clips. (2) We collect a user-rated Action-Design Scoring dataset comprising a total of 320 action sequences for eight intended messages and four representative eHMI modalities. The dataset validates that LLMs can translate intended messages into actions close to a human level, particularly for reasoning-enabled LLMs. (3) We introduce two automated raters, Action Reference Score (ARS) and Vision-Language Models (VLMs), to benchmark 18 LLMs, finding that the VLM aligns with human preferences yet varies across eHMI modalities.  [^1] The source code, prompts, Blender scenarios, and rendered clips are available at https://github.com/ApisXia/AutoActionDesign
@@
Ding Xia1, Xinyue Gui1, Fan Gao1, Dongyuan Li1 Mark Colley1,2, Takeo Igarashi1 1The University of Tokyo, 2University College London {dingxia1995, fangao0802, lidy94805}@gmail.com, gui-xinyue@g.ecc.u-tokyo.ac.jp, m.colley@ucl.ac.uk, takeo@acm.org
@@ 2.1 Rule-Based eHMI Action Planning
Currently, eHMI action planning generally follows a fixed design approach, in which human designers establish behavioral rules based on the specific features of different eHMI modalities. For example, in text- and icon-based eHMIs, designers create content referencing traffic regulation signs or standard messages Eisele and Petzoldt (2022); Eisma et al. (2021). In color- and light-band-based eHMIs, they design the content relying on human intuitive empathy and empirical evaluation with colors and blinking frequencies Bazilinskyy et al. (2019); Dey et al. (2020b). For anthropomorphic eHMIs, such as eyes or arms, designers mimic nonverbal communication cues drawn from common human-human interactions Mahadevan et al. (2018); Ochiai and Toyoshima (2011). Most recently, Colley et al. (2025) proposes using Human-In-The-Loop Multi-Objective Bayesian Optimization to create appropriate eHMIs. However, these eHMI action design only works as part of a case study to validate new eHMI modalities, which does not encourage the emergence of action design methods before our paper.
@@ 3.2 Action-Design Scoring Dataset
Robotic eyes are mounted on the front of the autonomous vehicle. The pupil’s position is specified using polar coordinates: the angle spans $0^{\circ},360^{\circ}]$ (starting from “up” and moving counterclockwise), and the distance spans $[0,1]$, where 0 denotes the center and $1$ is the edge Chang et al. (2022); Gui et al. (2022).
@@ 4.3 Benchmarking LLMs Performance
5 Discussion
Challenges in both the auto-design and rating processes. For LLM designers, we encountered two main problems. I) At the early stage of our prompt design, some models were “too lazy” to explore creative alternatives and copied the patterns of our examples, which are not always suitable. Encouraging more in-depth reasoning in the prompt helped mitigate this. II) We noticed that LLMs tend to include expressions of gratitude, but human designers prefer effectiveness, which made LLM-generated actions much longer than those written by humans. One possible remedy is to instruct the LLM to omit emotional expressions, but since emotion can be an essential part of some messages, finding the right balance between clarity and emotional tone – so that the actions feel human-like and satisfy human raters – remains a future direction.
For VLM raters, we identified two main challenges: differences from human annotators and limited recognition capability for small changes within an image series (Section 4.2). To address the first issue, we believe that collecting additional annotations from diverse human groups and fine-tuning VLMs to better align with human preferences could be effective. Regarding the second challenge, future improvements may involve utilizing more advanced VLM architectures.
Broader Implications and Applicability to Other Domains. Our method can be extended to domains such as social interaction, educational training, and caregiving Shiokawa et al. (2025), where the shared goal is to enable robots to perform actions that must be evaluated from a subjective perspective. For instance, indoor robots (e.g., vacuum cleaners) could use movements to convey alerting messages to homeowners in emergencies. Further, hand‐shaped robots might perform glove‐puppet shows to precisely convey the content of the story to children.
@@ 6 Conclusion
Acknowledgments
This work was supported by a Canon Research Fellowship and JST CRONOS, Grant Number JPMJCS24K8, Japan